# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 7.2. Generic KSBA stability of the mirror family
To prove ii), it is enough to prove that $Y_{t}$ is normal, since then the other conditions follow by i). By Theorem 7.5, the pair $(Y_{t},D_{t})$ is semi-log-canonical. In particular, the $Y_{t}$ is $S_{2}$, and so, by Serre’s criterion for normality [50, Thm 8.22A], it suffices to show that $Y_{t}$ is regular in codimension one. As in the proofs of Lemma 7.3 and Theorem 7.5, one can assume that there exists a good divisor $F$ on $\mathcal{X}$, inducing a $\mathbb{C}^{\star}$-action on $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$. The closure of the $\mathbb{C}^{\star}$-orbit passing through $t$ is an affine line $\mathbb{A}^{1}$, with $(\mathbb{A}^{1}\setminus\{0\})\subset T_{\mathcal{X}}$, and $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$, where $G=F^{\perp}\cap NE(\mathcal{X}/\overline{\mathcal{X}})$ and $J=\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]\setminus J$. The restriction $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ of $\mathcal{Y}_{\mathcal{X}}$ to this affine line is constant with fibers $Y_{t}$ on $\mathbb{A}^{1}\setminus\{0\}$. Moreover, by the proof of Lemma 7.2, the central fiber $\mathcal{Y}_{\mathcal{X},0}$ of $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ is a union of toric varieties with intersection complex $(P,\mathscr{P}_{J})$, where $\mathscr{P}_{J}$ is the subdivision of $P$ obtained from $\mathscr{P}_{J}$ by omitting the edges $e$ such that the corresponding curves $C_{e}$ in the double locus of $\mathcal{X}_{0}$ satisfy $F\cdot C_{e}=0$. In particular, the singular locus of $(P,\mathscr{P}_{J})$ is the union of its double curves $D_{e}$, which are in one-to-one correspondence with the edges $e$ of $\mathscr{P}_{J}$, that is, with the double curves $C_{e}$ of $\mathcal{X}_{0}$ such that $F\cdot C_{e}\neq 0$. Hence to prove that $Y_{t}$ is smooth in codimension one, it is enough to show that each of the double curves $D_{e}$ are generically smoothed out in $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$.
@@
It follows from the construction of $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ from the canonical scattering diagram and from [44, Eq. (2.17)] that a neighborhood of $D_{e}$ in $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ is isomorphic to the affine variety of equation $xy=t^{[C_{e}]}f_{W,e}$, where $f_{W,e}$ is the function of the canonical scattering diagram attached to the wall supported on the edge $e$ of $P$. As $(\mathbb{A}^{1}\setminus\{0\})$ is contained in $T_{\mathcal{X}}$, we have $t^{[C_{e}]}\neq 0$ on the non-central fibers of $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$. Moreover, by Lemma 6.9, the function $f_{W,e}$ is generically non-zero on $D_{e}$. This shows that all the double curves $D_{e}$ of $\mathcal{Y}_{\mathcal{X},0}$ are generically smoothed out in the one-parameter family $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$, and this concludes the proof that $Y_{t}$ is normal. ∎
@@ 7.3. KSBA stability of the mirror family
In this section, we prove the KSBA stability of the mirror family $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow S^{\mathrm{sec}}_{\mathcal{X}}$. To do this, we first provide an inductive description of the fibers.
@@
First note that $\mathcal{X}_{\sigma,0}$ is quasi-projective, reduced and normal crossings, as an open subset of the quasi-projective, reduced and normal crossings surface $\mathcal{X}_{0}$. If $v\in\sigma_{\mathbb{Z}}$ and $v\notin\partial\sigma$, then it follows from Lemma 7.8 ii) that $X^{v}_{\sigma}=X^{v}$, and so $X^{v}_{\sigma}$ is a 2-surface. If $v\in\partial\sigma$, then it follows from Definitions 3.2-3.4 that $X^{v}_{\sigma}$ is naturally a Zariski open subset of a 0-surface or a 1-surface if $\dim f_{H}(X^{v}_{\sigma})=2$ or $\dim f_{H}(X^{v}_{\sigma})=1$ respectively. Finally, it follows from [85, Proposition 3.3] that $\sigma$ is homeomorphic to a disk. ∎
Let $\mathcal{X}^{\prime}_{\sigma,0}$ be the quasi-projective open Kulikov surface given by Lemma 7.9. By Theorem 3.21, one can assume up to locally trivial deformations that $\mathcal{X}_{\sigma,0}^{\prime}$ is generic $d$-semistable, and so that $\mathcal{X}_{\sigma,0}^{\prime}$ is the central fiber of a quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$ by Theorem 3.26 and Lemma 3.31. The inclusions $\mathcal{X}_{\sigma,0}\subset\mathcal{X}_{0}$ and $\overline{\mathcal{X}}_{\sigma,0}\subset\mathcal{X}_{H,0}$ induce an injective map $NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})\rightarrow H=NE(\mathcal{X}/\mathcal{X}_{H})$, and so a dominant toric morphism
@@
The tropicalization of a punctured log curve $C$ contributing to $N_{pqr}^{\beta}(\mathcal{X})$ as in Definition 5.3 is a connected tropical curve in $C(P)$, with two unbounded edges with asymptotic directions $\mathbb{R}_{\geq 0}p$ and $\mathbb{R}_{\geq 0}q$, and with one leg contained in $\mathbb{R}_{\geq 0}r$. In particular, if $p,q,r$ were not all contained in a common cone $C(\sigma)$, this tropical curve would have to intersect a cone $C(\sigma\cap\sigma^{\prime})$ over the intersection of two distinct 2-dimensional faces $\sigma$, $\sigma^{\prime}$ of $\mathscr{P}_{\sigma}$, and so a cone $C(e)$ over an edge $e$ of $\mathscr{P}_{\sigma}$. In particular, the curve $C$ would contain an irreducible component with class a multiple of $[C_{e}]$, where $C_{e}$ is the double curve of $\mathcal{X}_{0}$ corresponding to $e$. Since $e$ is an edge of $\mathscr{P}_{\sigma}$, we have $[C_{e}]\notin H$, and this would imply $\beta\notin H$. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and let $H$ be a face of $NE(\mathcal{X}/\overline{\mathcal{X}})$. Then, for every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, for every $p,q,r\in C(\sigma)_{\mathbb{Z}}$ and $\beta\in H$, we have $N_{pqr}^{\beta}(\mathcal{X})=0$ if $\beta\notin NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$, and
$N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\beta}(\mathcal{X}_{\sigma})$
for every $\beta\in NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$.
@@
    The irreducible components $\mathcal{Y}_{\sigma}$ of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$ are in one-to-one correspondence with the 2-dimensional cells $\sigma$ of $\mathscr{P}_{H}$.
@@
    Denoting by $\partial\mathcal{Y}_{\sigma}$ the double locus of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$, and by $\mathcal{D}_{\sigma}$ and $\mathcal{C}_{\sigma}$ the restrictions of $\mathcal{D}_{\mathcal{X}}|_{S_{\mathcal{X}}}$ and $\mathcal{C}_{\mathcal{X}}|_{S_{\mathcal{X}}}$ to $\mathcal{Y}_{\sigma}$, the family $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})\rightarrow S_{H}$ is the base change, along the natural map $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ as in Equation (7.1), of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\sigma}})$ to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$.
@@
By Lemma 7.11, for every $p,q,r\in C(\sigma)_{\mathbb{Z}}$ and $\beta\in H$, we have $N_{pqr}^{\beta}(\mathcal{X})=0$ if $\beta\notin NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$, and $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\beta}(\mathcal{X}_{\sigma})$ for every $\beta\in NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$. Hence, $\mathcal{Y}_{\sigma}$ is the base change of $\mathcal{Y}_{\mathcal{X}_{\sigma}}$ along $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$, compatibly with the theta functions. In particular, it follows from the above description of $\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}$ in terms of theta functions that $\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}$ is the base change of $\mathcal{D}_{\mathcal{X}_{\sigma}}$, and, by Equation (6.6), $\mathcal{C}_{\sigma}:=\mathcal{C}_{\mathcal{X}}|_{\mathcal{Y}_{\sigma}}$ is the base change of $\mathcal{C}_{\mathcal{X}_{\sigma}}$.
Finally, the map $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ in Equation (7.1) is a dominant toric morphism and so restricts to a morphism $T_{H}\rightarrow T_{\mathcal{X}_{\sigma}}$. By Theorem 7.7 ii) applied to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$, the restriction $\mathcal{Y}_{\mathcal{X}_{\sigma}}|_{T_{\mathcal{X}_{\sigma}}}$ is irreducible. Hence, the restriction $\mathcal{Y}_{\sigma}|_{T_{H}}$ is also irreducible, and one concludes that $\mathcal{Y}_{\sigma}$ is an irreducible component of $\mathcal{Y}$. ∎
@@
Let $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ be a quasi-projective open Kulikov degeneration. Then, $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow S^{\mathrm{sec}}_{\mathcal{X}}$ is a family of KSBA stable pairs.
@@
By Theorem 7.6, we already have that $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}})\longrightarrow S^{\mathrm{sec}}_{\mathcal{X}}$ is a projective flat family of semi-log-canonical surfaces $(Y_{t},D_{t})$ such that $K_{Y_{t}}+D_{t}=0$. In particular, the divisor $K_{Y_{t}}+D_{t}+\epsilon\,C_{t}=\epsilon\,C_{t}$ is ample, and so it only remains to prove that the pairs $(Y_{t},D_{t}+\epsilon\,C_{t})$ are semi-log-canonical. It is enough to prove it over each of the torus orbits covering the toric variety $S^{\mathrm{sec}}_{\mathcal{X}}$.
@@
For a torus orbit corresponding to a cone of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$, one can assume, up to replacing $\mathcal{X}$ by another projective crepant resolution of $\overline{\mathcal{X}}$, that this cone is a face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$. By the description of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$ given in Theorem 7.12 i), and by [49, Proposition 7.2], it is enough to show that for every 2-dimensional face $\sigma$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$, the restriction of $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})$ to $T_{H}$ is a family of KSBA stable pairs. As the map $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ in Equation (7.1) is a dominant toric morphism, it restricts to a morphism $T_{H}\rightarrow T_{\mathcal{X}_{\sigma}}$. Hence, by Theorem 7.12 ii), the restriction of $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})$ to $T_{H}$ is a base change of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{X_{\sigma}})$ restricted to $T_{\mathcal{X}_{\sigma}}$. By Theorem 7.7 ii) applied to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$, the later family is a family of KSBA stable pairs and this concludes the proof.
@@
For a general torus orbit, one can assume as in Remark 6.29 that the corresponding cone is a bogus cone of the form $H+\sum_{v\in I}\mathbb{R}_{\geq 0}[X^{v}]$, with $H$ as above and $I$ a subset $P_{\mathbb{Z}}$ consisting of integral points $v$ such that $X^{v}$ is contracted by $f_{H}$. By Theorem 6.28, $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ in restriction to this torus orbit is obtained by $T_{0}$-equivariance from the restriction of $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ to $T_{H}$. As the base change in Theorem 7.12 is compatible with the $T_{0}$-actions, it is enough to show that, for every 2-dimensional face $\sigma$ of $\mathscr{P}_{H}$, the extension of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\sigma}})$ to the toric stratum corresponding to the bogus cone $\sum_{v\in I_{\sigma}}\mathbb{R}_{\geq 0}[X^{v}_{\sigma}]$ of $\mathcal{X}_{\sigma}$, where $I_{\sigma}=I\cap\sigma_{\mathbb{Z}}$, is a family of KSBA stable pairs. By Theorem 6.28, any fiber $Y_{\sigma,t}$ of this family is related by torus action, and so is isomorphic, to a fiber of $\mathcal{Y}_{\mathcal{X}_{\sigma}}|_{T_{\mathcal{X}_{\sigma}}}$, and so is normal by Theorem 7.7 ii). The KSBA stability then follows from Theorem 7.7 i). ∎

### 30 Mar 2024 v2

@@ 7.2. Generic KSBA stability of the mirror family
To prove ii), it is enough to prove that $Y_{t}$ is normal, since then the other conditions follow by i). By Theorem 7.5, the pair $(Y_{t},D_{t})$ is semi-log-canonical. In particular, the $Y_{t}$ is $S_{2}$, and so, by Serre’s criterion for normality [51, Thm 8.22A], it suffices to show that $Y_{t}$ is regular in codimension one. As in the proofs of Lemma 7.3 and Theorem 7.5, we can assume that there exists a good divisor $F$ on $\mathcal{X}$, inducing a $\mathbb{C}^{\star}$-action on $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$. The closure of the $\mathbb{C}^{\star}$-orbit passing through $t$ is an affine line $\mathbb{A}^{1}$, with $(\mathbb{A}^{1}\setminus\{0\})\subset T_{\mathcal{X}}$, and $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$, where $G=F^{\perp}\cap NE(\mathcal{X}/\overline{\mathcal{X}})$ and $J=\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]\setminus J$. The restriction $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ of $\mathcal{Y}_{\mathcal{X}}$ to this affine line is constant with fibers $Y_{t}$ on $\mathbb{A}^{1}\setminus\{0\}$. Moreover, by the proof of Lemma 7.2, the central fiber $\mathcal{Y}_{\mathcal{X},0}$ of $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ is a union of toric varieties with intersection complex $(P,\mathscr{P}_{J})$, where $\mathscr{P}_{J}$ is the subdivision of $P$ obtained from $\mathscr{P}_{J}$ by omitting the edges $e$ such that the corresponding curves $C_{e}$ in the double locus of $\mathcal{X}_{0}$ satisfy $F\cdot C_{e}=0$. In particular, the singular locus of $(P,\mathscr{P}_{J})$ is the union of its double curves $D_{e}$, which are in one-to-one correspondence with the edges $e$ of $\mathscr{P}_{J}$, that is, with the double curves $C_{e}$ of $\mathcal{X}_{0}$ such that $F\cdot C_{e}\neq 0$. Hence to prove that $Y_{t}$ is smooth in codimension one, it is enough to show that each of the double curves $D_{e}$ are generically smoothed out in $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$.
@@
It follows from the construction of $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ from the canonical scattering diagram, together with [45, Eq.(2.17)], that a neighborhood of $D_{e}$ in $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ is isomorphic to the affine variety of equation $xy=t^{[C_{e}]}f_{W,e}$, where $f_{W,e}$ is the function of the canonical scattering diagram attached to the wall supported on the edge $e$ of $P$. As $(\mathbb{A}^{1}\setminus\{0\})$ is contained in $T_{\mathcal{X}}$, we have $t^{[C_{e}]}\neq 0$ on the non-central fibers of $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$. Moreover, by Lemma 6.9, the function $f_{W,e}$ is generically non-zero on $D_{e}$. This shows that all the double curves $D_{e}$ of $\mathcal{Y}_{\mathcal{X},0}$ are generically smoothed out in the one-parameter family $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$, and this concludes the proof that $Y_{t}$ is normal. ∎
@@ 7.3. KSBA stability of the mirror family
In this section, we prove the KSBA stability of the mirror family $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$. To do this, we first provide an inductive description of the fibers.
@@
First note that $\mathcal{X}_{\sigma,0}$ is quasi-projective, reduced and normal crossings, as an open subset of the quasi-projective, reduced and normal crossings surface $\mathcal{X}_{0}$. If $v\in\sigma_{\mathbb{Z}}$ and $v\notin\partial\sigma$, then it follows from Lemma 7.8 ii) that $X^{v}_{\sigma}=X^{v}$, and so $X^{v}_{\sigma}$ is a 2-surface. If $v\in\partial\sigma$, then it follows from Definitions 3.2-3.4 that $X^{v}_{\sigma}$ is naturally a Zariski open subset of a 0-surface or a 1-surface if $\dim f_{H}(X^{v}_{\sigma})=2$ or $\dim f_{H}(X^{v}_{\sigma})=1$ respectively. Finally, it follows from [86, Proposition 3.3] that $\sigma$ is homeomorphic to a disk. ∎
Let $\mathcal{X}^{\prime}_{\sigma,0}$ be the quasi-projective open Kulikov surface given by Lemma 7.9. By Theorem 3.21, we can assume up to locally trivial deformations that $\mathcal{X}_{\sigma,0}^{\prime}$ is generic $d$-semistable, and so that $\mathcal{X}_{\sigma,0}^{\prime}$ is the central fiber of a quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$ by Theorem 3.26 and Lemma 3.31. The inclusions $\mathcal{X}_{\sigma,0}\subset\mathcal{X}_{0}$ and $\overline{\mathcal{X}}_{\sigma,0}\subset\mathcal{X}_{H,0}$ induce an injective map $NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})\rightarrow H=NE(\mathcal{X}/\mathcal{X}_{H})$, and so a dominant toric morphism
@@
The tropicalization of a punctured log curve $C$ contributing to $N_{pqr}^{\beta}(\mathcal{X})$ as in Definition 5.3 is a connected tropical curve in $C(P)$, with two unbounded edges with asymptotic directions $\mathbb{R}_{\geq 0}p$ and $\mathbb{R}_{\geq 0}q$, and with one leg contained in $\mathbb{R}_{\geq 0}r$. In particular, if $p,q,r$ were not all contained in a common cone $C(\sigma)$, this tropical curve would have to intersect a cone $C(\sigma\cap\sigma^{\prime})$ over the intersection of two distinct 2-dimensional faces $\sigma$, $\sigma^{\prime}$ of $\mathscr{P}_{\sigma}$, and so a cone $C(e)$ over an edge $e$ of $\mathscr{P}_{\sigma}$. In particular, the curve $C$ would contain an irreducible component with class a multiple of $[C_{e}]$, where $C_{e}$ is the double curve of $\mathcal{X}_{0}$ corresponding to $e$. Since $e$ is an edge of $\mathscr{P}_{\sigma}$, we have $[C_{e}]\notin H$, and this would imply $\beta\notin H$. Hence, the results follows since we assume that $\beta\in H$. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and let $H$ be a face of $NE(\mathcal{X}/\overline{\mathcal{X}})$. Then, for every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, for every $p,q,r\in C(\sigma)_{\mathbb{Z}}$ and $\beta\in H$, we have $N_{pqr}^{\beta}(\mathcal{X})=0$ if $\beta\notin NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$, and $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\beta}(\mathcal{X}_{\sigma})$ for every $\beta\in NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$.
@@
    There is a natural one-to-one correspondence between the irreducible components of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$ and the 2-dimensional cells $\sigma$ of $\mathscr{P}_{H}$.
@@
    For every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, denoting by $\mathcal{Y}_{\sigma}$ the corresponding irreducible component of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$, by $\partial\mathcal{Y}_{\sigma}$, $\mathcal{D}_{\sigma}$ and $\mathcal{C}_{\sigma}$ the restrictions of the double locus of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$, of $\mathcal{D}_{\mathcal{X}}|_{S_{\mathcal{X}}}$ and $\mathcal{C}_{\mathcal{X}}|_{S_{\mathcal{X}}}$ to $\mathcal{Y}_{\sigma}$, the family $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})\rightarrow S_{H}$ is the base change, along the natural map $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ as in Equation (7.1), of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\sigma}})$ to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$.
@@
By Lemma 7.11, for every $p,q,r\in C(\sigma)_{\mathbb{Z}}$ and $\beta\in H$, we have $N_{pqr}^{\beta}(\mathcal{X})=0$ if $\beta\notin NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$, and $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\beta}(\mathcal{X}_{\sigma})$ for every $\beta\in NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$. Hence, $\mathcal{Y}_{\sigma}$ is the base change of $\mathcal{Y}_{\mathcal{X}_{\sigma}}$ along $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$, compatibly with the theta functions. In particular, it follows from the above description of $\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}$ in terms of theta functions that $\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}$ is the base change of $\mathcal{D}_{\mathcal{X}_{\sigma}}$, and, by Equation (6.6), $\mathcal{C}_{\sigma}:=\mathcal{C}_{\mathcal{X}}|_{\mathcal{Y}_{\sigma}}$ is the base change of $\mathcal{C}_{\mathcal{X}_{\sigma}}$ along $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$.
Finally, the map $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ in Equation (7.1) is a dominant toric morphism and so restricts to a morphism $T_{H}\rightarrow T_{\mathcal{X}_{\sigma}}$. By Theorem 7.7 ii) applied to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$, the restriction $\mathcal{Y}_{\mathcal{X}_{\sigma}}|_{T_{\mathcal{X}_{\sigma}}}$ is irreducible. Hence, the restriction $\mathcal{Y}_{\sigma}|_{T_{H}}$ is also irreducible, and we conclude that $\mathcal{Y}_{\sigma}$ is an irreducible component of $\mathcal{Y}$. ∎
@@
Let $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ be a quasi-projective open Kulikov degeneration. Then, $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is a family of KSBA stable pairs.
@@
By Theorem 7.6, we already have that $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}})\longrightarrow S^{\mathrm{sec}}_{\mathcal{X}}$ is a projective flat family of semi-log-canonical surfaces $(Y_{t},D_{t})$ such that $K_{Y_{t}}+D_{t}=0$. In particular, the divisor $K_{Y_{t}}+D_{t}+\epsilon\,C_{t}=\epsilon\,C_{t}$ is ample, and so it only remains to prove that the pairs $(Y_{t},D_{t}+\epsilon\,C_{t})$ are semi-log-canonical. It is enough to prove this over each of the torus orbits covering the toric stack $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$.
@@
For a torus orbit corresponding to a cone of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$, we can assume, up to replacing $\mathcal{X}$ by another projective crepant resolution of $\overline{\mathcal{X}}$, that this cone is a face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$. By the description of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$ given in Theorem 7.12 i), and by [50, Proposition 7.2], it is enough to show that for every 2-dimensional face $\sigma$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$, the restriction of $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})$ to $T_{H}$ is a family of KSBA stable pairs. As the map $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ in Equation (7.1) is a dominant toric morphism, it restricts to a morphism $T_{H}\rightarrow T_{\mathcal{X}_{\sigma}}$. Hence, by Theorem 7.12 ii), the restriction of $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})$ to $T_{H}$ is a base change of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{X_{\sigma}})$ restricted to $T_{\mathcal{X}_{\sigma}}$. By Theorem 7.7 ii) applied to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$, the later family is a family of KSBA stable pairs and this concludes the proof in this case.
@@
For a general torus orbit, we can assume as in Remark 6.29 that the corresponding cone is a bogus cone of the form $H+\sum_{v\in I}\mathbb{R}_{\geq 0}[X^{v}]$, with $H$ as above and $I$ a subset $P_{\mathbb{Z}}$ consisting of integral points $v$ such that $X^{v}$ is contracted by $f_{H}$. By Theorem 6.28, the restriction of $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ to such a torus orbit is obtained by $T_{0}$-equivariance from the restriction of $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ to $T_{H}$. As the base change in Theorem 7.12 is compatible with the $T_{0}$-actions, it is enough to show that, for every 2-dimensional face $\sigma$ of $\mathscr{P}_{H}$, the extension of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\sigma}})$ to the toric stratum corresponding to the bogus cone $\sum_{v\in I_{\sigma}}\mathbb{R}_{\geq 0}[X^{v}_{\sigma}]$ of $\mathcal{X}_{\sigma}$, where $I_{\sigma}=I\cap\sigma_{\mathbb{Z}}$, is a family of KSBA stable pairs. By Theorem 6.28, any fiber $Y_{\sigma,t}$ of this family is related by torus action, and so is isomorphic, to a fiber of $\mathcal{Y}_{\mathcal{X}_{\sigma}}|_{T_{\mathcal{X}_{\sigma}}}$. By Theorem 7.7 ii), such a fiber is normal, and so the KSBA stability follows by Theorem 7.7 i). ∎